# Tendon-Based Proprioception in an Anthropomorphic Underactuated Robotic Hand with Series Elastic Actuators

### v1

@@ V Discussion & Conclusion
Despite these results, several limitations remain. First, the modeling neglects friction. While its effect was minimized in our system by using bearings and bushings, hardware with higher friction may experience larger estimation errors. Second, the approach depends on appropriate design parameters; for instance, overly weak torsional springs at the joints may lead to tension variations too small for reliable proximal contact detection. Finally, the interpretation of tension–excursion profiles is currently implemented through a rule-based algorithm, limiting robustness under unexpected or extreme grasping conditions. Future work will explore learning-based methods to process tension profile data, which may enable more generalized controllers capable of handling a wider range of situations.

### v2

@@ VI Discussion & Conclusion
Despite these results, several limitations remain. First, the estimation framework can be sensitive to friction as it is not explicitly modeled. While its effect was minimized in our design by using bearings and bushings, other hardware with higher friction may experience larger estimation errors. Second, the approach depends on appropriate design parameters; for instance, overly weak torsional springs at the joints may lead to tension variations too small for reliable proximal contact detection. Finally, the interpretation of tension–excursion profiles is currently implemented through a rule-based algorithm, limiting robustness under unexpected or extreme grasping conditions. Future work will explore learning-based methods to process tension profile, which may enable more generalized controllers capable of handling a wider range of situations.